# Pathfinder for Low-altitude Aircraft with Binary Neural Network

### v3

@@ III-B Binary Encoder for Feature Extraction
Camera Encoder: The camera encoder is a multi-scale binarized CNN-ViT hybrid model. We apply a shallow feature extraction network using convolutional modules to minimize computational complexity, while a deeper network utilizing ViT block focuses on capturing comprehensive features. Meanwhile, this approach addresses the problem of significant information loss associated with binarizing multiscale feature extraction frameworks. The image encoder operates through multiple downsampling modules that halve the size of the feature map at each layer, which helps to preserve critical details and reduces the loss of essential features. The CNN module applies a variant of ASPP named Multi-scale Binary dilated convolution Block (MBB, Fig.7) for the shallow network component. Meanwhile, this module retains substantial global and local information without increasing computational complexity, thereby enhancing overall performance.

### v4

@@ III-B Binary Encoder for Feature Extraction
Camera Encoder: The camera encoder is a multi-scale binarized CNN-ViT hybrid [45]. A shallow convolutional network (with Multi-scale Binary dilated convolution Block, MBB, Fig.7) reduces computation, while a deeper ViT block captures rich features. This design alleviates information loss from binarizing multi-scale frameworks. The encoder uses downsampling at each layer to preserve essential details and maintain overall performance.
@@ V CONCLUSIONS
ACKNOWLEDGEMENTS
This work was supported by the Fundo para o Desenvolvimento das Ciências e da Tecnologia of Macau (FDCT) with Reference No. 0067/2023/AFJ, No. 0117/2024/RIB2.